OpenStudy (anonymous):

If the perimeter of a rectangular property is 46 meters, and the area of the property is 76 meters, what is the length of each of the shorter sides

OpenStudy (anonymous):

Use 2(l+b)=46 and l*b=76 now solve for l

OpenStudy (anonymous):

ok now tell me who are you in your photo if your the girl you get help if boy ....you would still get help lol : )

OpenStudy (angela210793):

P=2x+2y=46--->x+y=23--->x=23-y A=x*y=76 x*y=76--->(23-y)y=76--->23y-y^2-76=0 y=4 or y=19 replace ys in the eq

OpenStudy (anonymous):

l +b =23 lb = 76 l^2 + 76 - 23l = 0 l = 19 b =4 Now tell me who are you ?

OpenStudy (anonymous):

guy

OpenStudy (anonymous):

me too welcome to guy club hehe lol